RHIC Startup:
Get ready to run 78x78 with reduced deuteron intensity. The TSPs were
activated yesterday during a ramp, so it is important to see if this brought
down the background rates in STAR.
One of the studies last night was to prepare 3 fills with d intensity of 1.0, 0.8 and 0.6 x 10^11 ions. (Keith was able to inject 1.4E11 for one store) But ya1 had tripped off. Ron Schroeder was providing a noisy BBQ signal to increase the Blue emittance, from 14-15 to 17-18. Comparisons of the lifetime of the three different intensity bunches show that over a twenty minute store, there was no difference between the three different bunch sets.
The BtA foil drive is not functioning well. The current foil cannot be moved. We ran last night with the Au foil left in for D running. Keith reports that he can't tell the difference in the D beam. It may give a slightly increased emittance in Blue, which is the direction we want to move anyway. It seems as if running until the next maintenance day (2 weeks) without fixing it is good. The drive is binding, but can still be moved by hand. Continuing to move it will cause it to seize. .
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |